A pen holder holds twelve pens, five of which do not write. If five pens are chosen at random, find the probability that at least one
(a) writes. (b) does not write.
Please Please include full step by step, I have the answer, I need to see the work for my test, I just can't figure it out and can't find another example of this. THANK YOU!

Answer by sudhanshu_kmr(1152)


number of ways to select 5 pens out of 12= 12C5 = 792
probability that no one write P(A)= 5C5/792 = 1/792
a) probability that at least one write P(A')= 1 - P(A) = 791/792

probability that all write P(B) = 7C5/792 = 21/792
b)
probability that at least does not write P(B')= 1 - P(B) = 771/792


if any doubt, you are welcome to contact.